Grade 11Mechanics

Q)A CAR IS TRAVELLING ON A STRAIGHT ROAD.THE MAX VELOCITY THE CAR CAN ATTAIJN IS 24 M/S.THE MAX ACCELERATION AND RETARDATION IT CAN ATTAIN IS 1 AND 4 RESPECTIVELY.THE SHORTEST TIME THE CAR TAKES FROM REST TO REST IN A DISTANCE OF 200M IS....

ANS)22.4S

sam 94
15 Years agoGrade 11

1 Answer

AJIT AskiitiansExpert-IITD
ApprovedApproved Tutor Answer15 Years ago

Dear Koka Ram ,

 Let us assume that upto distance x it accelerates and so 200-x it deaccelerates.and let the peak velocity be v.

   V = 1*t1    -->  t1 = v

  X = (1/2) * 1* t1^2  

  x = v^2/2        ----- (1)

  now for deaccleration ,

  0 - V = -4*t2  -----> t2 = v/4

  200 -x = v * v/4 - (1/2)*4*(v/4)^2     ------(2)

add (1) and (2)

200 = v2 /2 + v/ 4 - v2 /8

200 = 5v2 /8

v = 17.88

 

therefore total time = v +v/4 =  22.4 s

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE

preparation.

 

All the best.

 

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query

on askiitians forum and become an Elite Expert League askiitian.

 

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar

respectively : Click here to download the toolbar..

 

Askiitians Expert

 

Ajit Singh Verma

IIT Delhi